I am a parent who trying to help her son and not doing to well. One of the problems is 1 3/5 x 2 1/2. It comes out to 4/1, but tdoes that mean it then would a whole 4 or 4/1?
Thank you Answer by ptaylor(2198) 
 4/1 is the same as 4 because 4 divided by 1=4
1 3/5 times 2 1/2
1 3/5=8/5
and
2 1/2=5/2
(8/5)(5/2)=the 5's cancel so we have left 8/2=4
Hope this helps and good luck----ptaylor